Reserves Liabilities $19,200 Deposits $240,000 220.800 Loans Refer to Table 1. Assume the Fed's new reserve requirement is 6 percent and that the Bank of Springfield makes new loans so as to make their required reserves just 6 percent of their deposits , From then on no bank holds any excess reserves . Assume also that people hold only deposits and no currency . By what amount does the economy's money supply increase because of the the new loans? Select one: 2. \$80,000 O b \$50,200 $72,000 d. \$106,000
Answer #1

Answer:

Given:

Reserve requirement = 6% = .06

Money multiplier = 1/0.06 = 16.67

Reserve requirement = 6 % of Deposits = 6% of 240,000 = 14400

Excess Reserve = Reserves Liabilities - Reserve requirement = 19200 - 14400 = 4800

Change in money supply = Change in reserves x money multiplier

Change money supply = 4800 x 16.67 = 80000

Money supply increases by 80000

Hence Correct option is A] 80,000